Which of the following is larger? Please explain how you made your decisions.

3 fifths or 0.59
6 fourths or 1.6
2 over 1000 or 0.0025

Answer by Cintchr(481)

3 fifths or 0.59
3/5 = 6/10 = .6
3/5 is greater


6 fourths or 1.6
6/4 = 3/2 = 1.5
1.6 is greater


2 over 1000 or 0.0025
2/1000 = 0.002 = 0.0020
0.0025 is greater